OpenStudy (anonymous):

Three of four nubers have a sum of 22, if the average of the four numbers is 8, what is the fourth number? How do I figure that out?

OpenStudy (anonymous):

Let the four numbers be \(w,x,y,z\). You know that the sum of three of these is 22. Let \(w,x,y\) be these three numbers. \[w+x+y=22\] You also know that the average of all four numbers is 8: \[\frac{w+x+y+z}{4}=8\] Substitute the first equation into this one: \[\frac{22+z}{4}=8\] Now you can solve for the fourth number, \(z\).

OpenStudy (anonymous):

I honestly have not taken a math class in 6 years, and have forgotten how to solve even  Please help?

OpenStudy (anonymous):

\[\frac{22+z}{4}=8\\ 4\cdot\frac{22+z}{4}=4\cdot8\\ 22+z=32\\ 22+z-22=32-22\\ ~~~~~~~~~~~~~~~~~~~\vdots\]

OpenStudy (anonymous):

so, z=10?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome!